The cubic equation z^3+Az^2+Bz+26=0, A, BeR(A, B are elements in the set R) one root of the above cubic equation is 1+i. Find the real root (ii) Determine the values of A and B.
1
Expert's answer
2020-11-09T20:19:37-0500

First of all, as coefficients A and B are real, we know that if "z \in \mathbb{C}" is a root of an equation, its' complex conjugate "\bar{z}" is also a root of this equation :

"\bar{z}^3+A\bar{z}^2+B\bar{z}+26=\bar{z^3}+\bar{(Az^2)}+\bar{(Bz)}+26=\bar{0}=0"

Therefore, "\bar{(1+i)} = 1-i" is also a root of the equation.

Now, to find the third root we can apply, for example, Vieta's formulas :

"z_1\times z_2 \times z_3 = (-1)^3 26"

"(1-i)\times(1+i)\times z_3 = -26"

"z_3 = -13"

To find the coefficients we can either replace "z" in the equation by the roots we've found (which is more direct method), either reapply the Vieta's formulas :

"A = -z_1-z_2-z_3=11"

"B = z_1z_2+z_2z_3+z_1z_3 = -24"

We can even replace "z" in our equation by its' 3 different values and verify that these calculations are correct.